OpenStudy (anonymous):

nth derivative of y = cos5x.cos3x

OpenStudy (shubhamsrg):

2cos a cos b = cos(a-b) + cos(a+b) thus cos5x cos3x = (1/2)(cos 2x + cos 8x) rest should be easier i guess?

OpenStudy (shubhamsrg):

i can post the answer if you want..but you should try..